OpenStudy (anonymous):

Find the distance between the points (6, 5) and (4, 3).

OpenStudy (anonymous):

10,8?

OpenStudy (anonymous):

\[\sqrt{\left( 6-4 \right)^{2}+\left( 5-3 \right)^{2}} = \sqrt{4+4} = \sqrt{8} = 2\sqrt{2}\]

OpenStudy (anonymous):

no sorry thats wrong